# Causal LLM Routing: End-to-End Regret Minimization from Observational Data

### v1

@@ Abstract
LLM routing aims to select the most appropriate model for each query, balancing competing performance metrics such as accuracy and cost across a pool of language models. Prior approaches typically adopt a decoupled strategy, where the metrics are first predicted and the model is then selected based on these estimates. This setup is prone to compounding errors and often relies on full-feedback data, where each query is evaluated by all candidate models, which is costly to obtain and maintain in practice. In contrast, we learn from observational data, which records only the outcome of the model actually deployed. We propose a causal end-to-end framework that learns routing policies by minimizing decision-making regret from observational data. To enable efficient optimization, we introduce two theoretically grounded surrogate objectives: a classification-based upper bound, and a softmax-weighted regret approximation shown to recover the optimal policy at convergence. We further extend our framework to handle heterogeneous cost preferences via an interval-conditioned architecture. Experiments on public benchmarks show that our method outperforms existing baselines, achieving state-of-the-art performance across different embedding models.
@@ 1 Introduction
LLM routing is an emerging research area focused on optimizing model selection for each input query, balancing performance and cost across a pool of available LLMs. Because LLM performance varies significantly by task and input (Hu et al., 2024), as well as computational cost (Ong et al., 2024), dynamic routing strategies have been proposed to select the most suitable model per query (Shnitzer et al., 2023). This challenge becomes even more critical in agentic applications, where multiple LLM calls may be made within a single workflow, making efficient model selection essential for user experience and resource allocation. As LLM deployment scales, routing also contributes to environmental sustainability by reducing unnecessary computation (Singh et al., 2025).
Routing methods can be broadly classified based on whether they invoke one or multiple models per query. Multi-model approaches include non-predictive routing, which cascades models sequentially from light to heavy (Wang et al., 2023; Chen et al., 2023), and predictive ensembles, which learn to combine outputs or scores from multiple LLMs (Shekhar et al., 2024; Huang et al., 2025). While these methods can improve accuracy and robustness, they incur significant computational cost and latency due to repeated inference. By contrast, predictive routing methods, including our approach, select a single LLM for each query by training a router that maps input queries to the most appropriate model (Shnitzer et al., 2023; Ong et al., 2024; Somerstep et al., 2025). This framework offers a more scalable and cost-efficient solution, particularly in settings where minimizing latency and compute resources is critical.
@@
Another fundamental limitation in the existing literature on predictive routing is its reliance on full-feedback datasets. Prior work (Ong et al., 2024) (Somerstep et al., 2025) assumes access to data where each query has been evaluated by all available LLMs. This assumption is impractical: (i) the computational and monetary cost of exhaustively querying all models is prohibitive; and (ii) the rapid pace of LLM development makes it challenging to maintain comprehensive, up-to-date evaluation datasets. In contrast, observational data, where each query is evaluated by only one model, is readily available from real-world LLM deployments, making it far more scalable than full feedback datasets. However, it introduces treatment bias from historical routing policies, which can lead to suboptimal decisions if not properly addressed (Swaminathan and Joachims, 2015; Künzel et al., 2019).
To the best of our knowledge, this is the first work to (i) learn LLM routing from observational data and (ii) introduce an integrated learning framework on routing. Our main contributions are:
@@ 2.2 End-to-End Regret Minimization
Our goal is to route each prompt $x$ to the LLM $t$ such that the decision leads to the highest possible utility $y_{x}(t)$. To this end, we aim to learn an end-to-end policy $f:\mathcal{X}\to\mathcal{T}$ that minimizes the decision-making regret (Fernández-Loría and Provost, 2022; Zou et al., 2022). More formally,
@@
$Regret(f):=\mathbb{E}_{X}[Y_{X}(t^{*})-Y_{X}(f(X))]=\mathbb{E}_{X}[Y_{X}(t^{*})]-\mathbb{E}_{X}[Y_{X}(f(X))],$ (1)
with $t^{*}:=\arg\max_{t\in\mathcal{T}}Y_{X}(t)$.
@@
2.2.1 Surrogate Loss 1: Classification-Based Upper Bound
Our first approach is to derive a tractable upper bound on the regret and directly minimize it. To do so, we introduce the following regularity assumption:
Assumption 1 ($L$-Lipschitz Continuity).
We assume that the estimated utility function $\hat{Y}_{x}(\cdot)$ is $L$-Lipschitz continuous with constant $L>0$ under distance metric $d:\mathcal{T}\times\mathcal{T}\to\mathbb{R}_{\geq 0}$.
This assumption is mild and standard in structured decision-making settings. Since $\mathcal{T}$ is a finite set and $\hat{Y}_{x}(\cdot)$ is typically learned via smooth function approximators (e.g., neural networks), it is natural to expect bounded variation in utility values across nearby treatments.
@@ 2.2.1 Surrogate Loss 1: Classification-Based Upper Bound
Under Assumption 1, the regret in Equation (2) can be upper bounded by
$\text{Regret}(f)\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}d(t_{i}^{*},f(x_{i})),\quad\text{where }t_{i}^{*}:=\arg\max_{t\in\mathcal{T}}\hat{Y}_{x_{i}}(t).$ (3)
@@ 2.3 Estimating Counterfactual Utility via Causal Inference
In the previous sections, we assumed that we have access to $\hat{Y}_{X}(\cdot)$. Given the observational nature of the data, the potential utility function $Y_{x}(\cdot)$ is not directly observable. We follow the potential outcomes framework Rosenbaum and Rubin (1983); Rubin (1984) and assume the existence of a potential utility function $Y_{x}(t)$. We adopt the following assumptions, which are standard in the causal inference literature.
@@
The assigned treatments and potential outcomes are independent conditional on observed covariates, i.e. $t\perp\{Y_{x}(t^{\prime})|t^{\prime}\in\mathcal{T}\}|x$ (Hirano and Imbens, 2004).
@@
In the causal inference literature, counterfactual outcomes can be estimated using methods such “meta-learner” (Künzel et al., 2019), or the Inverse Propensity Weighting (IPW) estimator (Horvitz and Thompson, 1952). In this work, we utilize the doubly robust estimator introduced by Dudík et al. (2011), defined as:
@@
The doubly robust estimator combines an outcome model $\hat{r}_{t}(x)$ and a propensity model $\hat{p}(t|x)$, yielding consistent estimates if either is correctly specified (Dudík et al., 2011). It offers a favorable bias-variance trade-off: the propensity model corrects for selection bias, while the outcome model reduces variance by leveraging structure in the data.
@@ 3.1 Interval-Conditioned Joint Router
We assume access to a finite, representative set of cost preferences $\Lambda:=\{\lambda_{1},\dots,\lambda_{m}\}\subset\mathbb{R}_{\geq 0}$. For ease of notation we assume that $\lambda_{1}<\lambda_{2}<\dots<\lambda_{m}$. In practice, these values may correspond to discrete service quality tiers (e.g., basic, standard, premium) that reflect users’ varying willingness to trade off cost for performance. For each $\lambda\in\Lambda$, we partition the training data by cost preference and estimate $\lambda$-specific utility $\hat{Y}_{x_{i}}^{\lambda}(t)$, which forms the basis for our joint interval-conditioned architecture.
@@
    For each interval $[\lambda_{j},\lambda_{j+1}]\,j\in[1,\dots,m-1]$, we initialize a joint network $f(x,\lambda):\mathcal{X}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{|\mathcal{T}|}$ that uses as input both $x$, $\lambda\in(\lambda_{j},\lambda_{j+1})$.
@@ 3.2 Model Architecture
A key component of the heterogeneous cost preference routing setup introduced in Section 3.1 is the interval-conditioned joint model $f(x,\lambda)$, that for each interval $[\lambda_{j},\lambda_{j+1}]$, interpolates between the two corresponding individual models $f_{\lambda_{j}}$ and $f_{\lambda_{j+1}}$. Specifically, the architecture is designed to exploit the affine structure of the utility function with respect to $\lambda$, namely $y=a-\lambda c$. This motivates a lightweight parameterization that uses only the two endpoints of the interval $[\lambda_{j},\lambda_{j+1}]$ rather than all $m$ pre-trained models. Concretely, the joint model that we propose is defined as:
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation, and $g(\lambda):=\texttt{Activation}(\texttt{Linear}(\lambda))$ is a learnable representation of the cost sensitivity parameter.
This architecture enables smooth interpolation between $f_{\lambda_{j}}$ and $f_{\lambda_{j+1}}$ within $[\lambda_{j},\lambda_{j+1}]$, allowing the router to adapt to intermediate values of $\lambda$ without requiring an individual model for each one. By conditioning only on the two bounding models, this design achieves computational efficiency and strong generalization across cost preferences. The proposed architecture is illustrated in Figure 1.
@@ 4.1 Datasets
We evaluate our methods on two publicly available benchmarks for LLM routing: RouterBench (Hu et al., 2024) and SPROUT (Somerstep et al., 2025).
@@
SPROUT is a larger and more diverse benchmark focused on cost-aware routing, consisting of 44,241 prompts and responses from 13 state-of-the-art LLMs. The prompts cover six challenging tasks: GPQA (Rein et al., 2024), MuSR (Sprague et al., 2023), MMLU-Pro (Wang et al., 2024), MATH (Hendrycks et al., 2021), OpenHermes (Teknium, 2023), and RAGBench (Friel et al., 2024). SPROUT includes a predefined split: 80% for training, with the remaining 20% evenly divided between validation and test sets.
@@ 4.2 Embeddings and Model Architecture
To encode input queries into vector representations $x$, we generate embeddings using two compact, publicly available language models: BERT-base-uncased (768 dimensions) and Llama-3.2-1B (2048 dimensions). Each input is passed once through the model, and the final hidden states are mean-pooled to obtain a fixed-length embedding. These models were selected for their efficiency and suitability for real-time routing.
The embeddings are processed by a two-layer fully connected network with GELU activations and 200 hidden units per layer. The model is trained with the Adam optimizer (learning rate $1\times 10^{-4}$) for up to 10,000 epochs, using early stopping with a patience of 100. A softmax output with temperature $\tau=100$ is used to control the sharpness of the output probabilities. This architecture is used consistently across all benchmarked methods for fair comparison. For the doubly robust estimator, the same network models the direct outcomes $\hat{r}_{t}(x)$, while the propensity scores $\hat{p}(t\mid x)$ are estimated using XGBoost. To reduce variance from extreme inverse propensity weights, we apply clipping at the 5th and 95th percentiles. The only architectural modification is for the interval-based model, where the softmax temperature is increased to $\tau=1000$ to enable smoother interpolation across $\lambda$. Hyperparameters are summarized in Appendix C.
@@ 4.3 Methods
As an optimistic oracle, we include a Full-Feedback model that learns a model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{T}|}$ using the complete outcome vector for each query and optimizes a standard multi-class classification loss. We benchmark against a common decoupled routing strategy denoted Baseline, which independently estimates model accuracy $a_{x}(t)$ and cost $c_{x}(t)$, without accounting for selection bias. This reflects the approach taken in prior predictive routing methods such as CARROT (Somerstep et al., 2025), which has demonstrated superior performance over alternatives like RouteLLM (Ong et al., 2024) and RoRF (Jain et al., 2023). To adjust for treatment assignment bias, we consider a Regress-and-Compare (R&C) method, which fits outcome models $\hat{Y}_{x}(t)$ for each treatment $t$, and selects the action $\hat{t}=\arg\max_{t}\hat{Y}_{x}(t)$. Building on this, we implement a Causal-CARROT variant by adapting both the parametric and kNN instantiations of CARROT to the R&C framework. We additionally include CF-Regression, which models $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{T}|}$ and is trained to minimize MSE against the counterfactual utility function from the doubly robust estimator: $\min_{f}\ \sum_{i=1}^{n}\sum_{t=1}^{|\mathcal{T}|}(\hat{Y}_{x_{i}}(t)-f(x_{i})_{t})^{2}.$ Decisions are made by selecting the treatment with the highest predicted value, i.e., $\hat{t}=\arg\max_{t}f(x)_{t}$.

### v2

@@ Abstract
LLM routing aims to select the most appropriate model for each query, balancing competing performance metrics such as accuracy and cost across a pool of language models. Prior approaches typically adopt a decoupled strategy, where metrics are first predicted and the model is then selected based on these estimates. This setup is prone to compounding errors and often relies on full-feedback data, where each query is evaluated by all candidate models, which is costly to obtain and maintain in practice. In contrast, we learn from observational data, which records only the outcome of the model actually deployed. We propose a causal end-to-end framework that learns routing policies by minimizing decision-making regret from observational data. To enable efficient optimization, we introduce two theoretically grounded surrogate objectives: a classification-based upper bound, and a softmax-weighted regret approximation shown to recover the optimal policy at convergence. We further extend our framework to handle heterogeneous cost preferences via an interval-conditioned architecture. Experiments on public benchmarks show that our method outperforms existing baselines, achieving state-of-the-art performance across different embedding models.
@@ 1 Introduction
LLM routing is an emerging research area focused on optimizing model selection for each input query, balancing performance and cost across a pool of available LLMs. Because LLM performance varies significantly by task and input (hu2024routerbench), as well as computational cost (ong2406routellm), dynamic routing strategies have been proposed to select the most suitable model per query (shnitzer2023large). This challenge becomes even more critical in agentic applications, where multiple LLM calls may be made within a single workflow, making efficient model selection essential for user experience and resource allocation. As LLM deployment scales, routing also contributes to environmental sustainability by reducing unnecessary computation (singh2025survey).
Routing methods can be broadly classified based on whether they invoke one or multiple models per query. Multi-model approaches include non-predictive routing, which cascades models sequentially from light to heavy (wang2023fusing; chen2023frugalgpt), and predictive ensembles, which learn to combine outputs or scores from multiple LLMs (shekhar2024towards; huang2025thriftllm). While these methods can improve accuracy and robustness, they incur significant computational cost and latency due to repeated inference. By contrast, predictive routing methods, including our approach, select a single LLM for each query by training a router that maps input queries to the most appropriate model (shnitzer2023large; ong2406routellm; somerstep2025carrot). This framework offers a more scalable and cost-efficient solution, particularly in settings where minimizing latency and compute resources is critical.
@@
Another fundamental limitation in the existing literature on predictive routing is its reliance on full-feedback datasets. Prior work (ong2406routellm) (somerstep2025carrot) assumes access to data where each query has been evaluated by all available LLMs. This assumption is impractical: (i) the computational and monetary cost of exhaustively querying all models is prohibitive; and (ii) the rapid pace of LLM development makes it challenging to maintain comprehensive, up-to-date evaluation datasets. In contrast, observational data, where each query is evaluated by only one model, is readily available from real-world LLM deployments, making it far more scalable than full feedback datasets. However, it introduces treatment bias from historical routing policies, which can lead to suboptimal decisions if not properly addressed (swaminathan2015batch; kunzel2019metalearners).
To the best of our knowledge, this is the first work to (i) learn LLM routing from observational data and (ii) introduce an integrated learning framework for routing. Our main contributions are:
@@ 2.2 End-to-End Regret Minimization
Our goal is to route each prompt $x$ to the LLM $t$ such that the decision leads to the highest possible utility $y_{x}(t)$. To this end, we aim to learn an end-to-end policy $f:\mathcal{X}\to\mathcal{T}$ that minimizes the decision-making regret (fernandez2022causal; pmlr-v162-zou22a). More formally,
@@
$Regret(f):=\mathbb{E}_{X}[Y_{X}(t^{*}_{X})-Y_{X}(f(X))]=\mathbb{E}_{X}[Y_{X}(t^{*}_{X})]-\mathbb{E}_{X}[Y_{X}(f(X))],$ (1)
with $t^{*}_{X}:=\arg\max_{t\in\mathcal{T}}Y_{X}(t)$.
@@
Surrogate Loss 1: Classification-Based Upper Bound
Our first approach is to derive a tractable upper bound on the regret and directly minimize it. To do so, we define the following notion of Lipschitz continuity for utility functions over the probability simplex.
Definition 1.
Let $\hat{Y}_{x}:\mathcal{T}\to\mathbb{R}$ be an estimated utility function assigning a scalar utility to each model $t\in\mathcal{T}$ for a given input $x$. We extend $\hat{Y}_{x}$ to the probability simplex $\Delta^{|\mathcal{T}|}$ by defining
$\hat{Y}_{x}(p):=\sum_{t\in\mathcal{T}}p(t)\hat{Y}_{x}(t),$ (3)
for any $p\in\Delta^{|\mathcal{T}|}$. We say that $\hat{Y}_{x}$ is $L$-Lipschitz over the simplex with respect to the $\ell_{1}$ norm if for all $p,q\in\Delta^{|\mathcal{T}|}$,
$|\hat{Y}_{x}(p)-\hat{Y}_{x}(q)|\leq L\cdot\|p-q\|_{1}.$ (4)
Here, the constant $L$ can be taken as $L:=\max_{t\in\mathcal{T}}|\hat{Y}_{x}(t)|$.
This condition ensures that small changes in the model distribution lead to bounded changes in expected utility. Since $\mathcal{T}$ is a finite set and $\hat{Y}_{x}(\cdot)$ is typically learned via bounded smooth function approximators (e.g., neural networks), it is natural to expect bounded variation in utility values across nearby treatments. Note that when $p$ is a one-hot vector $e_{t^{*}}$ and $q=f(x)$ is a probabilistic policy, this setting corresponds to our model selection problem, where we seek to minimize the regret between the optimal choice and a stochastic routing decision.
@@
Suppose the estimated utility function $\hat{Y}_{x}:\mathcal{T}\to\mathbb{R}$ is $L$-Lipschitz continuous over the probability simplex with respect to the $\ell_{1}$ norm, as in Definition 1. Then, for a policy $f:\mathcal{X}\to\Delta^{|\mathcal{T}|}$ that outputs a distribution $f(x)$ over $\mathcal{T}$, the regret can be upper bounded by:
$\text{Regret}(f)\leq L\cdot\frac{1}{n}\sum_{i=1}^{n}\sqrt{2\cdot\text{CE}(t_{i}^{*},f(x_{i}))},$ (5)
where $t_{i}^{*}:=\arg\max_{t\in\mathcal{T}}\hat{Y}_{x_{i}}(t)$ is the optimal treatment for input $x_{i}$, and $\text{CE}(t_{i}^{*},f(x_{i})):=-\log f(x_{i})_{t_{i}^{*}}$ denotes the cross-entropy loss.
@@ 2.3 Estimating Counterfactual Utility via Causal Inference
In the previous sections, we assumed that we have access to $\hat{Y}_{X}(\cdot)$. Given the observational nature of the data, the potential utility function $Y_{x}(\cdot)$ is not directly observable. We follow the potential outcomes framework rosenbaum1983central; rubin1984bayesianly and assume the existence of a potential utility function $Y_{x}(t)$. We adopt the following assumptions, which are standard in the causal inference literature.
@@
The assigned treatments and potential outcomes are independent conditional on observed covariates, i.e. $t\perp\{Y_{x}(t^{\prime})|t^{\prime}\in\mathcal{T}\}|x$ (hirano2004propensity).
@@
In the causal inference literature, counterfactual outcomes can be estimated using various methods. Such examples include the “meta-learner” (kunzel2019metalearners), or the Inverse Propensity Weighting (IPW) estimator (horvitz1952generalization). In this work, we utilize the doubly robust estimator introduced by dudik2011doubly, defined as:
@@
The doubly robust estimator combines an outcome model $\hat{r}_{t}(x)$ and a propensity model $\hat{p}(t|x)$, yielding consistent estimates if either is correctly specified (dudik2011doubly). It offers a favorable bias-variance trade-off: the propensity model corrects for selection bias, while the outcome model reduces variance by leveraging structure in the data.
Remark 1.
While we use the doubly robust (DR) estimator in our experiments due to its favorable bias–variance tradeoff and strong practical performance, our framework is estimator-agnostic: any valid counterfactual estimator can be used to compute utility estimates. This includes more advanced approaches that relax or mitigate these assumptions
@@ 3.1 Interval-Conditioned Joint Router
We assume access to a finite, representative set of cost preferences $\Lambda:=\{\lambda_{1},\dots,\lambda_{m}\}\subset\mathbb{R}_{\geq 0}$. For ease of notation we assume that $\lambda_{1}<\lambda_{2}<\dots<\lambda_{m}$. In practice, these values may correspond to discrete service quality tiers (e.g., basic, standard, premium) that reflect users’ varying willingness to trade off cost for performance. For each $\lambda\in\Lambda$, we partition the training data by cost preference and estimate $\lambda$-specific utility $\hat{Y}_{x_{i}}^{\lambda}(t)$, which forms the basis of our joint interval-conditioned architecture.
@@
    For each interval $\lambda_{j},\lambda_{j+1}]\,j\in[1,\dots,m-1]$, we initialize a joint network $f(x,\lambda):\mathcal{X}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{|\mathcal{T}|}$ that uses as input both $x$, $\lambda\in(\lambda_{j},\lambda_{j+1})$.
@@ 3.2 Model Architecture
A key component of the heterogeneous cost preference routing setup introduced in Section 3.1 is the interval-conditioned joint model $f(x,\lambda)$, that for each interval $\lambda_{j},\lambda_{j+1}]$, interpolates between the two corresponding individual models $f_{\lambda_{j}}$ and $f_{\lambda_{j+1}}$. Specifically, the architecture is designed to exploit the affine structure of the utility function with respect to $\lambda$, namely $y=a-\lambda c$. This motivates a lightweight parameterization that uses only the two endpoints of the interval $\lambda_{j},\lambda_{j+1}]$ rather than all $m$ pre-trained models. Concretely, the joint model that we propose is defined as:
@@
where $\cdot$$\cdot$] denotes concatenation, and $g(\lambda):=\texttt{Activation}(\texttt{Linear}(\lambda))$ is a learnable representation of the cost sensitivity parameter.
This architecture enables smooth interpolation between $f_{\lambda_{j}}$ and $f_{\lambda_{j+1}}$ within $\lambda_{j},\lambda_{j+1}]$, allowing the router to adapt to intermediate values of $\lambda$ without requiring an individual model for each one. By conditioning only on the two bounding models, this design achieves computational efficiency and strong generalization across cost preferences. The proposed architecture is illustrated in Figure 1.
@@ 4.1 Datasets
We evaluate our methods on two publicly available benchmarks for LLM routing: RouterBench (hu2024routerbench) and SPROUT (somerstep2025carrot).
@@
SPROUT is a larger and more diverse benchmark focused on cost-aware routing, consisting of 44,241 prompts and responses from 13 state-of-the-art LLMs. The prompts cover six challenging tasks: GPQA (rein2024gpqa), MuSR (sprague2023musr), MMLU-Pro (wang2024mmlu), MATH (hendrycks2021measuring), OpenHermes (teknium2023openhermes), and RAGBench (friel2407ragbench). SPROUT includes a predefined split: 80% for training, with the remaining 20% evenly divided between validation and test sets.
@@ 4.2 Embeddings and Model Architecture
To encode input queries into vector representations $x$, we generate embeddings using two compact, publicly available language models: BERT-base-uncased (768 dimensions) and Llama-3.2-1B (2048 dimensions). Each input is passed once through the model, and the final hidden states are mean-pooled to obtain a fixed-length embeddings. These models were selected for their efficiency and suitability for real-time routing.
The embeddings are processed by a two-layer fully connected network with GELU activations and 200 hidden units per layer. The model is trained with the Adam optimizer (learning rate $1\times 10^{-4}$) for up to 10,000 epochs, using early stopping with a patience of 100. A softmax output with temperature $\tau=100$ is used to control the sharpness of the output probabilities. This architecture is used consistently across all benchmarked methods for fair comparison. For the doubly robust estimator, the same network models the direct outcomes $\hat{r}_{t}(x)$, while the propensity scores $\hat{p}(t\mid x)$ are estimated using XGBoost. To reduce variance from extreme inverse propensity weights, we apply clipping at the 5th and 95th percentiles. The only architectural modification is for the interval-based model, where the softmax temperature is increased to $\tau=1000$ to enable smoother interpolation across $\lambda$. Hyperparameters are summarized in Appendix D.
@@ 4.3 Methods
As an optimistic oracle, we include a Full-Feedback model that learns a model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{T}|}$ using the complete outcome vector for each query and optimizes a standard multi-class classification loss. We benchmark against a common decoupled routing strategy denoted Baseline, which independently estimates model accuracy $a_{x}(t)$ and cost $c_{x}(t)$, without accounting for selection bias. This reflects the approach taken in prior predictive routing methods such as CARROT (somerstep2025carrot), which has demonstrated superior performance over alternatives like RouteLLM (ong2406routellm) and RoRF (jain2023rorf). To adjust for treatment assignment bias, we consider a Regress-and-Compare (R&C) method, which fits outcome models $\hat{Y}_{x}(t)$ for each treatment $t$, and selects the action $\hat{t}=\arg\max_{t}\hat{Y}_{x}(t)$. Building on this, we implement a Causal-CARROT variant by adapting both the parametric and kNN instantiations of CARROT to the R&C framework. We additionally include CF-Regression, which models $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{T}|}$ and is trained to minimize MSE against the counterfactual utility function from the doubly robust estimator: $\min_{f}\ \sum_{i=1}^{n}\sum_{t=1}^{|\mathcal{T}|}(\hat{Y}_{x_{i}}(t)-f(x_{i})_{t})^{2}.$ Decisions are made by selecting the treatment with the highest predicted value, i.e., $\hat{t}=\arg\max_{t}f(x)_{t}$.